If f(x)=3x^2+3x-3 and I need to solve for f(x+3) how would I do this step.
Answer by fractalier(6550)

Given this function
f(x)=3x^2+3x-3, we can find
f(x+3) by plugging in x+3 for every x we see in the original...we get
f(x+3) = 3(x+3)^2 + 3(x+3) - 3 = 3(x^2+6x+9) + 3x+9 - 3 =
3x^2 + 18x + 27 + 3x + 9 - 3 =
3x^2 + 21x + 33